How much load is needed to elongate a steel wire to 0.5 %of its original length ? The diameter of the wire is 2mm and youngest modulus of the material of the wire is 20×10^10 N/m^2

1
Expert's answer
2022-01-28T08:14:15-0500

Given:

"\varepsilon=0.005"

"r=1\:\rm mm=10^{-3}\:\rm m"

"E=20*10^{10}\:\rm N/m^2"


The Hooke's law says

"\sigma=\frac{F}{A}=\varepsilon E"

Hence, the load

"F=\varepsilon AE=\pi\varepsilon r^2E"

"F=3.14*0.005*(10^{-3})^2*20*10^{10}=\rm 3140\: N"